# Nested and outlier embeddings into trees

### v1

@@ Abstract
In this paper, we consider outlier embeddings into HSTs and ultrametrics. In particular, for $(X,d)$, let $k$ be the size of the smallest subset of $X$ such that all but that subset (i.e. the “outlier set”) can be probabilistically embedded into the space of HSTs with expected distortion at most $c$. Our primary result is showing that there exists an efficient algorithm that takes in $(X,d)$ and a target distortion $c$ and samples from a probabilistic embedding with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and distortion at most $(32+\epsilon)c$, for any $\epsilon>0$. This leads to better instance-specific approximations for certain instances of the buy-at-bulk and dial-a-ride problems, whose current best approximation algorithms go through HST embeddings.
In order to facilitate our results, we largely focus on the concept of compositions of nested embeddings introduced by [20]. A nested embedding is a composition of two embeddings of a metric space $(X,d)$ – a low distortion embedding of a subset $S$ of nodes, and a higher distortion embedding of the entire metric. The composition is a single embedding that preserves the low distortion over $S$ and does not increase distortion over the remaining points by much. In this paper, we expand this concept from the setting of deterministic embeddings to the setting of probabilistic embeddings. We show how to find good nested compositions of embeddings into HSTs, and combine this with an approximation algorithm of Munagala et al. [39] to obtain our results.
@@ 1 Introduction
Low distortion metric embeddings are an indispensable algorithmic tool in many contexts. Many algorithmic problems are easier to solve on trees than on general metrics, so embedding a given metric into a tree and solving the problem on the tree provides a way to obtain an approximately optimal solution on the original metric. However, the final approximation factor depends on the distortion (the multiplicatively worst change in any pairwise distance) of the embedding. This motivates the problem of finding an embedding from general metrics into trees of the least possible distortion. A series of works [3, 11, 12, 28] showed that although some metrics require linear distortion to embed into trees, probabilistic embeddings can do much better: it is possible to embed any $n$ point metric into a distribution over trees such that the expected distance between any pair of points increases by at most $O(\log n)$ [28], while the cycle requires a factor of $\Omega(n)$ for any deterministic tree embedding [3]. In fact, the probabilistic upper bound applies to embeddings into a special family of trees called hierarchically separated trees, or HSTs. The particular structure of these rooted trees – with edge lengths decreasing geometrically over levels going from the root to a leaf – is algorithmically useful in a variety of problems.
While uniform distortion bounds are helpful in understanding the general effectiveness of using metric embeddings to solve a problem, from a practical perspective, a more pertinent question is how well a particular given metric $(X,d)$ can embed into trees. Indeed, for some metrics the optimal distortion is $\Omega(\log n)$ but for others it may be much smaller, e.g. $O(1)$, providing much better algorithmic results. Is it possible to efficiently obtain the best embedding for the given metric into trees? Munagala et al. [39] considered this question from the viewpoint of approximation and provided a constant factor approximation to the instance-optimal distortion.
In this paper we further investigate instance-optimal distortion for probabilistic embeddings into trees by exploring the role of outliers. For a finite metric $(X,d)$, a $(k,c)$-outlier embedding is a map from “most of” $X$, namely all but $k$ vertices that we call outliers, into another metric space $Y$ such that the distortion over the non-outliers is bounded by $c$. Intuitively, we can view these outliers as noisy or adversarial data that may hurt the quality of the instance-optimal embedding from $X$ into $Y$. The goal of the outlier embedding is to identify and remove these outliers upfront so as to obtain a much better embedding over the remaining points. Algorithmically, given a target distortion $c$, we want to find the smallest $k$ such that the removal of $k$ outliers enables embedding the rest of the metric with distortion at most cc. [^1] In Section 3.1, we will discuss some specific examples where removing a small number of points can greatly improve overall distortion. For deterministic embeddings, this problem turns out to be NP-hard when the host metric is a tree or ultrametric [44]. We extend the notion of outlier embeddings to probabilistic output, and give bicriteria approximations focusing on minimizing the outlier set size when given a target distortion for the final probabilistic embedding. Our main result is as follows.
Theorem 1.1:
@@
This result allows us to obtain improved instance-specific approximation ratios for several graph optimization problems (e.g. the buy-at-bulk and dial-a-ride problems) on input metrics that meet certain structural requirements. We describe these applications in Section 3.
@@
A key technical tool in our approach is the concept of compositions of nested embeddings that was first introduced by Chawla and Sheridan [20] in the context of outlier embeddings into the Euclidean metric. A nested embedding seeks to take some “good” embedding on a subset $S\subseteq X$, as well as a set of “okay” embeddings on all subsets of $X$ with size at most $|X\setminus S|+1$; It then finds a new embedding on the entire set $X$ such that distortion is still good for pairs of points in $S$ and is not too bad for any pair of points. Chawla and Sheridan use nested embeddings to ensure a bound on the optimal solution to an SDP, which they then round to determine their outlier set and embedding.
We combine the ideas of Chawla and Sheridan with the approximation algorithm of Munagala et al. [39] to obtain outlier embeddings into HSTs. Munagala et al.’s algorithm rounds a natural linear programming (LP) relaxation of probabilistic embeddings into HSTs. We extend this LP to incorporate outliers; develop a nested composition into HSTs to guide the choice of distortion constraints in the LP; and provide a rounding similar to that in [39]. The way we use nested compositions to enable finding bicriteria approximations through an LP rounding scheme uses some of the techniques of Chawla and Sheridan [20]. However, we still have two key technical challenges.
Nested compositions into HSTs with bounded contraction.
To bound the modified LP of Munagala et al. [39] that we use in this setting, we must show existence of nested compositions into HSTs that are non-contracting. In particular, Munagala et al.’s LP essentially sets up a metric labeling problem with the goal of assigning each vertex to a representative at geometrically increasing levels. Each vertex $v$ is assigned a representative at each level $i$ that is at a distance no more than $2^{i}$ from $v$. Importantly, the level-wise labeling constraints cannot be weakened for fractional outliers – each vertex, whether or not it’s an outlier, must still be assigned at level $i$ to a representative within distance $2^{i}$ from it. Even slightly weakening this constraint can potentially introduce errors at all levels of the hierarchy, leading to very large distortion (depending on $|X|$ rather than $k$ and $c$).
Nested compositions successively “stitch together” pairs of embeddings on the outliers and the non-outliers. For normed metrics, such as those considered by [20], when contraction is allowed, it suffices to directly concatenate the respective embeddings. Our setting, in contrast, calls for a careful alignment and stitching of the two embeddings that does not contract any distances. We call this procedure a “merge” and characterize the conditions under which it returns a correct nested embedding. We show that such a merge function exists for nested embeddings into HSTs.
@@
The second technical challenge is in extending the notion of nested embeddings to probabilistic embeddings. [^2] Note that while the embeddings are probabilistic, we require the outlier set to be deterministic. In particular, our nested embedding guarantees are designed specifically with respect to a deterministic outlier set as, by definition, we obtain tighter bounds on the distortion on non-outliers relative to the outliers. Allowing the outlier set to be random is an interesting direction for future work, but would require a significant modification to the manner in which we use nested embeddings to obtain outlier embeddings. Although we are not aware of explicit examples of metrics where removing a random set of outliers provides even better distortion than removing a deterministic outlier set, it is possible that this extension would provide further improvement in the instance-specific approximation factors for some applications. This is important in the context of tree embeddings because, as mentioned previously, probabilistic tree embeddings can obtain exponentially smaller distortion relative to deterministic tree embeddings.
Consider extending a probabilistic embedding of some subset $S$ of a given metric $(X,d)$ into a host space $Y$ to a probabilistic embedding of the entire metric $X$ into $Y$. One potential way of leveraging a deterministic nested embedding is to draw a random embedding $\alpha$ mapping $S$ to $Y$, extend this embedding, and return the extension. What can we say about the distortion of this random extension? Let $c_{\alpha}:=\max_{u,v\in X}(d_{\alpha}(u,v)/d(u,v))$ denote the (probabilistic) distortion achieved by the (probabilistic) embedding $\alpha$, where $d_{\alpha}$ is the distance function in the host metric composed with the mapping $\alpha$. Then, the deterministic nested embedding we construct guarantees us a distortion linear in $c_{\alpha}$. This implies that the distortion of the final random embedding returned by this approach is at worst linear in $\mathbb{E}_{\alpha}[c_{\alpha}]$. However, $\mathbb{E}_{\alpha}[c_{\alpha}]$, can be much larger than the expected distortion of the original embedding, $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$. To avoid this issue, we show that our nested embeddings into HSTs satisfy a certain obliviousness property, and we show that this property lets us obtain a final distortion guarantee that depends on $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$.
To summarize, our primary contributions are as follows.
  • We extend the concept of nested embeddings to the setting of probabilistic embeddings. We give a general algorithm for composing probabilistic nested embeddings into a given host space, so long as we can design a “merge” algorithm with certain properties. We show that the set of embeddings into HSTs has such a merge algorithm and as a result has such an algorithm for composing embeddings.
  • We discuss how to extend the notion of outlier sets to probabilistic embeddings and show that if a given metric $(X,d)$ has an embedding into HSTs with at most $c$ distortion and at most $k$ outliers, we can sample from a probabilistic embedding with at most $(32+\epsilon)c$ distortion and at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers, for any $\epsilon>0$.
  • We also achieve similar bicriteria approximation results for the weighted outlier problem, where the cost of an outlier set is the sum of the weights assigned to its nodes, and the goal is to find a minimum cost outlier set such that the remaining metric embeds into HSTs with distortion $c$.
  • We present several algorithmic applications of outlier embeddings into HSTs.
@@
Embeddings are a widely studied concept, and there are many algorithms that obtain “uniform” distortion bounds for embedding arbitrary $n$-point metrics into a particular target space. A series of results [3, 12, 18, 28] beginning with Bartal [11] studied how to probabilistically embed a general metric into HSTs with only poly-logarithmic distortion. The seminal paper of Fakcharoenphol, Rao, and Talwar [28] showed that all $n$-point metrics have a probabilistic embedding into the space of hierarchically separated trees with distortion at most $O(\log n)$. We will use this general algorithm to construct embeddings on small subsets of a metric that we hope to compose with an “optimal” embedding on some larger subset of that metric.
@@
There is also a variety of papers that study instance-optimal distortion for embedding finite metrics into a given target space. That is, the goal in such embeddings is to minimize the distortion needed for the given specific $(X,d)$, which in some cases may be smaller than the best bound achievable for all metrics of that size. Some examples of such papers cover embeddings into constant dimensional Euclidean space [8, 26, 27, 37, 45], the line [37, 29, 7, 41], trees [21, 9], and ultrametrics [4, 39].
Because of their simple structure, HST embeddings are useful for a wide variety of algorithmic applications, including ones related to the $k$-server problem [24, 10], hierarchical clusterings [15, 43, 25], Steiner trees [30], buy-at-bulk network design [6], metric labeling [34], vehicle routing problems [32], and online clustering [22].
@@
The notion of outlier embeddings was initially defined by Sidiropoulos, Wang, and Wang [44] where they studied deterministic embeddings into ultrametrics, trees, and constant-dimensional $\ell_{2}$ space. But their techniques are largely unrelated to ours. One part of their work considers deterministic, isometric outlier embeddings (where the distortion $c$ is 1), and uses structural properties to find approximations. These structural properties don’t apply to non-isometric embeddings. Another part considers bicriteria approximations, but bounds additive distortion (that is, the maximum absolute difference between old and new distances, rather than their ratio). Techniques for bounding additive distortion tend to be quite different from those for multiplicative distortion, the focus of this work. Moreover, additive distortion does not provide approximation guarantees for algorithmic applications.
@@
Chubarian and Sidiropoulos [23] continued the work on outlier embeddings and studied non-isometric outlier embeddings of unweighted graph metrics into the line. Later, Chawla and Sheridan [20] studied deterministic outlier embeddings into $\ell_{2}$ space with unbounded dimension. There is a notion related to outlier embeddings, called embeddings with slack, in which the goal is to minimize the number of edges that are distorted by more than some target factor, and there has been significant work in that area as well [1, 16, 35]. Outlier embeddings are also closely related to metric Ramsey theory [13, 38]. Given inputs $n$ and $c$, metric Ramsey theory asks: What is the largest $m$ such that all metrics of size $n$ have a submetric of size $m$ that embeds into some target space with distortion at most $c$? In this paper we are primarily focused on instance-optimal solutions rather than universal bounds that apply to all metrics, and we seek to approximate the size of the outlier set rather than the size of the non-outlier set. Existing results for metric Ramsey theory give universal bounds for deterministic embeddings into $\ell_{p}$ spaces and ultrametrics.
@@
Our notion of nested embeddings was first defined in [20], but many other types of embedding compositions or extensions have been considered before. One particular example is the notion of a Lipschitz extension [33, 40, 14], which seeks to leave non-outliers unaltered from the original embedding and ensure the expansion (but not necessarily the contraction) of every pair of nodes is bounded. In fact, our nested embedding algorithm is very similar to a part of the Lipschitz extension algorithm presented by Calinescu et al. [14] as a part of their solution to the 0-extension problem. Other works study a variety of different embedding compositions [2, 5, 19, 42, 36]. These notions largely differ from ours in that they emphasize bounding the worst distortion for any pair in the entire metric, while our notion focuses on bounding the worst distortion between any pair subject to a “special set” having even stricter distortion constraints.
@@ 2 Definitions and results
In this section, we will define terms used in this paper and summarize our results.
@@
A metric space is a set of points $X$ with distances $d_{X}(\cdot,\cdot)$ between them that obey the triangle inequality. That is, for all $x,y,z\in X$, $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$. We assume without loss of generality (via scaling) that in the metrics we consider, the minimum distance between distinct pairs of points is at least $1$: $\min_{x\neq y\in X}d_{X}(x,y)\geq 1$.
A deterministic embedding of a space $(X,d_{X})$ into a space $(Y,d_{Y})$ is a map $\alpha:X\rightarrow Y$. We are interested in the distortion (i.e. multiplicative change) of pairwise distances under this map. We use, as a convenient shorthand, $d_{\alpha}(x,y):=d_{Y}(\alpha(x),\alpha(y))$ to denote the distance between the images of two points $x,y\in X$ under this map, and $d(x,y)$ to denote the original distance $d_{X}(x,y)$ when the space $(X,d_{X})$ is clear from the context. We will without loss of generality consider only expanding embeddings, or embeddings for which $d_{\alpha}(x,y)\geq d(x,y)$ for all $x,y\in X$.
We can further expand these definitions to the setting of probabilistic embeddings. In particular, $\mathcal{D}$ is a probabilistic embedding of $(X,d_{X})$ into $\mathcal{Y}$ if for each function $\alpha$ in the support of $\mathcal{D}$, $\alpha:X\rightarrow Y(\alpha)$ for some $(Y(\alpha),d_{Y(\alpha)})\in\mathcal{Y}$. This embedding has distortion $c$ if for all $u,v\in X$ we have:
@@
$\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq c\cdot d(u,v).$
Observe that every embedding in the support of $\mathcal{D}$ is expanding, but pairwise distortion is bounded in expectation over the randomness in $\mathcal{D}$.
@@
In this paper, we extend the concept of outlier embeddings to its probabilistic counterpart. Note that probabilistic embeddings can, in general, obtain much smaller distortion into a class of metric spaces than deterministic embeddings, which motivates extending the study of such embeddings to this setting. If $\mathcal{D}$ is a probabilistic embedding from $(X,d_{X})$ into $\mathcal{Y}$ and $K\subseteq X$, define $\mathcal{D}_{X\setminus K}$ to be the distribution identical to $\mathcal{D}$ but with all functions $\alpha$ in the support restricted to the set $X\setminus K$ (i.e. we can sample from $\mathcal{D}_{X\setminus K}$ by sampling from $\mathcal{D}$ and restricting the domain of the function drawn).
Definition 2.1:
@@
Note that in Definition 2.1, there is a deterministic subset $K$ whose distances to other nodes (and each other) we don’t care about. The randomness is only on the embedding for the remaining nodes in $X\setminus K$. One motivation for outlier embeddings is as a way to prevent noise or adversarial data from preventing good embeddings, and we believe that making the outlier set itself deterministic is consistent with this motivation. In Section 3.4 we will also discuss a specific problem where the application naturally lends itself to finding a deterministic outlier set.
Composition of nested embeddings
Chawla and Sheridan [20] introduce the concept of a composition of nested embeddings. This concept considers two (partial) embeddings from a metric space $(X,d_{X})$ into space $(Y,d_{Y})$. One embedding $\alpha_{S}$ maps a subset $S\subset X$ into $Y$ with distortion $c_{S}$, and the other maps the entire space $X$ into $Y$ with distortion $c_{X}>c_{S}$. The goal is to combine these embeddings into a single embedding from $X$ into $Y$ that preserves the distortion $c_{S}$ over pairs vertices in $S$ and obtains distortion comparable to $c_{X}$ over the remaining pairs of vertices.
We will now extend this concept to a probabilistic setting. We assume that both the “good” embedding from $S$ into a space $\mathcal{Y}$ as well as the “poor” embedding from $X$ into $\mathcal{Y}$ are probabilistic. Furthermore, as in [20] the composition only needs to apply the “poor” embedding over subsets of $X$ of size at most $k+1$ where $k:=|X\setminus S|$. Therefore, for the tightest distortion bounds, we will assume that we are given an assortment of probabilistic embeddings (one for each subset of $X$ of size $k+1$) into $\mathcal{Y}$ and we will use $c_{X}$ to denote the worst distortion of any embedding in this assortment.
Definition 2.2 (Probabilistic composition of nested embeddings, modified from [20]):
@@
A (strong) $(\ell,g)$-nested composition of the assortment $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$, where $\ell$ is a constant and $g:[1,\infty)^{2}\times\mathbb{N}\rightarrow[1,\infty)$, is an algorithm that given the assortment returns a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ such that:
@@
$\text{for all }u,v\in S,$ $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq\ell\cdot c_{S}\cdot d(u,v),$ (2)
$\text{and, for all }u,v\in X,$ $\mathop{E}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq g(c_{S},c_{X},k)\cdot d(u,v).$ (3)
Chawla and Sheridan primarily considered weak nested compositions that do not require non-contraction of distances for pairs of points not both in $S$. For our application to outlier embeddings into HSTs, restricting contraction of all points is necessary.
We do consider weak nested compositions as well, and we will briefly discuss some minor results for probabilistic Lipschitz extensions in Appendix D.
@@
In this paper, we are particularly interested in embedding into distributions of hierarchically separated trees, which are metrics defined as follows:
Definition 2.3:
@@
  • Each node $u$ in the tree is associated with some value $\eta_{u}$, and $\eta_{v}=\beta\cdot\eta_{u}$ whenever $u$ is a child of $v$.
  • For all leaves $v$, $\eta_{v}=1/\beta$.  [^3] In this paper, we will use the term $\beta$-HSTs to refer to what other sources sometimes call exact $\beta$-HSTs. These sources typically define non-exact $\beta$-HSTs as only requiring $\eta_{v}\geq\beta\cdot\eta_{u}$.
  • For two leaves of the tree $x,y\in X$ with least common ancestor $z$, $d(x,y)=\eta_{z}$.
For simplicity, we will primarily focus on $2$-HSTs, and in referring to HSTs, we mean $2$-HSTs unless otherwise noted. Note that focusing on $\beta$-HSTs for a different constant $\beta$ at worse changes results by a constant factor. We may also write $\mathcal{H}$ to mean the set of $2$-HSTs.
@@
In order to achieve our main outlier embedding result, we will show in Section 4 that under the right conditions there exist good nested compositions of probabilistic embeddings. The right conditions will depend on the existence of a separate “merging” algorithm for combining embeddings, and we will show in Section 5 that such an algorithm exists for HSTs. This gives us following theorem that we will use to prove Theorem 2.5.
Theorem 2.4 (HSTs have strong nested embeddings):

### v2

@@ Abstract
In this paper, we consider outlier embeddings into HSTs. In particular, for metric $(X,d)$, let $k$ be the size of the smallest subset of $X$ such that all but that subset (the “outlier set”) can be probabilistically embedded into the space of HSTs with expected distortion at most $c$. Our primary result is showing that there exists an efficient algorithm that takes in $(X,d)$ and a target distortion $c$ and samples from a probabilistic embedding with at most $O(\frac{k}{\epsilon}\log^{2}k)$ outliers and distortion at most $(32+\epsilon)c$, for any $\epsilon>0$. In order to facilitate our results, we show how to find good nested embeddings into HSTs and combine this with an approximation algorithm of Munagala et al. [31] to obtain our results.
@@ 1 Introduction
Many algorithmic solutions to metric problems begin by embedding the given metric into a hierarchically separated tree (HST) and then solving the problem on that more structured space. The distortion (the multiplicatively worst change in pairwise distance) of the embedding affects the final approximation factor of such algorithms. Randomized algorithms that optimize the expected value of the objective function can use probabilistic embeddings, which bound the expected distortion for each pair of nodes. A series of works [4, 10, 11, 23] showed that probabilistic embeddings can embed any $n$-point metric into trees with expected distortion $O(\log n)$.
While uniform distortion bounds are helpful in understanding the general effectiveness of using metric embeddings to solve a problem, from a practical perspective, we may ask how well a particular given metric $(X,d)$ can embed into trees. Indeed, for some metrics the optimal distortion is $\Omega(\log n)$ but for others it may be much smaller, e.g. $O(1)$, providing better algorithmic results.
In this paper we investigate instance-optimal distortion for probabilistic embeddings into trees by exploring the role of outliers. For a finite metric $(X,d)$, a $(k,c)$-outlier embedding is a map from “most of” $X$, namely all but $k$ vertices that we call outliers, into another metric space $Y$ such that the distortion over the non-outliers is bounded by $c$. Intuitively, we can view the outliers as noisy or adversarial data that may hurt the quality of the instance-optimal embedding from $X$ into $Y$. The goal of the outlier embedding is to identify and remove these outliers upfront so as to obtain a much better embedding over the remaining points. Algorithmically, given a target distortion $c$, we want to find the smallest $k$ such that the removal of $k$ outliers enables embedding the rest of the metric with distortion at most $c$. In this paper, we focus on outlier embeddings in a probabilistic setting, and give bicriteria approximations that provide tradeoffs between outlier set size and distortion. Our main result is as follows.
Theorem 1.1
@@
Munagala et al. [31] provide an approximation algorithm for the instance-optimal probabilistic embedding into HSTs that rounds a natural linear programming (LP) relaxation. We extend this LP to incorporate outliers. Proving the feasibility of our extended LP requires the concept of nested embeddings, that was first developed by Chawla and Sheridan [19] in the context of deterministic embeddings into normed metrics. We extend this notion to probabilistic HST embeddings, thereby guiding the choice of distortion constraints in the LP; and then provide a rounding similar to that in [31].
Informally, a nested embedding takes as input an embedding on a subset $S\subseteq X$ with low distortion $c_{S}$, as well as a set of embeddings on all subsets of $X$ with size at most $|X\setminus S|+1$, each with (larger) distortion $c_{X}$; it then finds a single new embedding on the entire set $X$ such that distortion is close to $c_{S}$ for pairs of points in $S$ and is only a polylogarithmic factor worse than $c_{X}$ over other pairs. Extending Chawla and Sheridan [19]’s technique to HSTs poses two key technical challenges that we will now discuss.
Nested embeddings into HSTs with bounded contraction.
Chawla and Sheridan [19] approximate optimal outlier sets by rounding an SDP solution. The SDP has designated “outlier” variables $\delta_{i}$ such that when $\delta_{i}$ is set to $1$, the expansion and contraction conditions for the distance between $i$ and any other node are relaxed; in fact, the contraction can be unbounded. Thus, [19] need only show existence of nested embeddings that bound expansion, not contraction.
By contrast, while our modified version of the LP from Munagala et al. [31] also includes outlier variables $\delta_{i}$, a large $\delta_{i}$ does not admit contraction between nodes, so bounding the optimal LP solution requires nested embeddings that bound both expansion and contraction. In particular, Munagala et al.’s LP essentially sets up a metric labeling problem with the goal of assigning each vertex to a representative at geometrically increasing levels. Each vertex $v$ is assigned a representative at each level $i$ that is at a distance no more than $2^{i}$ from $v$. Importantly, the level-wise labeling constraints cannot be weakened for fractional outliers – each vertex, whether or not it’s an outlier, must still be assigned at level $i$ to a representative within distance $2^{i}$ from it. Even slightly weakening this constraint can potentially introduce errors at all levels of the hierarchy, leading to very large distortion (depending on $|X|$ rather than $k$ and $c$).
In contrast to the main outlier embedding results of Chawla and Sheridan [19], our setting calls for a careful alignment and stitching of the two embeddings that does not contract any distances. We call this procedure a “merge” and characterize the conditions under which it returns a correct nested embedding. Chawla and Sheridan implicitly give such a merge function for $\ell_{1}$ embeddings, and we show that such a merge function exists for HST embeddings.
@@
The second technical challenge is in extending the notion of nested embeddings to probabilistic embeddings. This is important in the context of tree embeddings because probabilistic tree embeddings can obtain exponentially smaller distortion than deterministic tree embeddings.
Consider taking a probabilistic embedding of some subset $S$ of a given metric $(X,d)$ into a host space $Y$ and extending it to a probabilistic embedding of the entire metric $X$ into $Y$. One potential way of leveraging a deterministic nested embedding is to draw a random embedding $\alpha$ mapping $S$ to $Y$, extend this embedding, and return the extension. What can we say about the distortion of this random extension? We are guaranteed distortion at most $\mathbb{E}_{\alpha}[\max_{u,v\in X}(d_{\alpha}(u,v)/d(u,v))]$, but this can be much larger than the expected distortion of the original embedding, $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$. To avoid this issue, we show that our nested embeddings into HSTs satisfy a certain obliviousness property, and that this property lets us obtain a final distortion guarantee that depends on $\max_{u,v\in X}(\mathbb{E}_{\alpha}[d_{\alpha}(u,v)]/d(u,v))$.
@@
Embedding into HSTs has been studied primarily with the objective of minimizing “uniform” distortion. See e.g. [4, 10, 11, 16, 23]. In our work, we will use as a subroutine the algorithm of Fakcharoenphol et al. [23], which for any $n$ point metric samples from an $O(\log n)$ distortion probabilistic embedding into HSTs. Instance-optimal distortion has been studied for target spaces such as constant dimensional Euclidean space [8, 13, 22, 29, 35], the line [29, 24, 7, 33], trees [20, 9], and ultrametrics [3, 31]. The goal in such embeddings is to minimize the distortion needed for the given specific $(X,d)$, which in some cases may be smaller than the best bound achievable for all metrics of that size.
@@
The notion of outlier embeddings was initially defined by Sidiropolous et al. [36] when they studied deterministic embeddings into ultrametrics, trees, and constant-dimensional $\ell_{2}$ space. They consider approximating deterministic, isometric outlier embeddings (where the distortion $c$ is 1) and bicriteria approximations with additive distortion. The structural properties used for isometric approximations don’t apply to non-isometric embeddings, and techniques for bounding additive distortion tend to be quite different from those for multiplicative distortion, the focus of this work. Moreover, additive distortion does not provide approximation guarantees for algorithmic applications.
Chubarian and Sidiropolous [21] continued the work on outlier embeddings and studied non-isometric outlier embeddings of unweighted graph metrics into the line. Later, Chawla and Sheridan [19] studied deterministic outlier embeddings into $\ell_{2}$ space with unbounded dimension. The techniques of [19] are similar to the those of Calinescu et al. [14] for the 0-extension problem. The primary differences are: (1) [14] use a semimetric obtained from an LP solution to the 0-extension problem everywhere that [19] use the input metric $d$, and (2) [14] pick a uniformly random permutation over the set $S$ of nodes in the good embedding (which they refer to as terminals), but [19] take a random permutation over the set of $X\setminus S$’s closest neighbors in $S$. We will follow the analysis of [19], but a modified version of the analysis in [14] provides another approach.
@@
Another related notion is embeddings with slack, in which the goal is to minimize the number of edges that are distorted by more than some target factor [1, 15, 27]. Outlier embeddings are also closely related to metric Ramsey theory [12, 30]. Given inputs $n$ and $c$, metric Ramsey theory asks: What is the largest $m$ such that all metrics of size $n$ have a submetric of size $m$ that embeds into some target space with distortion at most $c$? In this paper we are primarily focused on instance-optimal solutions rather than universal bounds that apply to all metrics, and we seek to approximate the size of the outlier set rather than the size of the non-outlier set.
@@
Our notion of nested embeddings first appears in [19], but many other types of embedding compositions or extensions have been considered. One particular example (which we discuss further in Appendix 0.C) is the notion of a Lipschitz extension [26, 32, 14], which seeks to leave non-outliers unaltered from the original embedding and ensure the expansion (but not necessarily the contraction) of every pair of nodes is bounded. Other works study a variety of different embedding compositions [2, 5, 17, 34, 28]. These notions largely differ from ours in that they emphasize bounding the worst distortion for any pair in the entire metric, while our notion focuses on bounding the worst distortion between any pair subject to a “special set” having even stricter distortion constraints.
@@ 2 Definitions and results
A metric space is a set of points $X$ with distances $d_{X}(\cdot,\cdot)$ between them that obey the triangle inequality. That is, for all $x,y,z\in X$, $d_{X}(x,z)\leq d_{X}(x,y)+d_{X}(y,z)$. We assume without loss of generality (via scaling) that in the metrics we consider, the minimum distance between distinct pairs of points is at least $1$. If $\alpha:X\rightarrow Y$ is a map from $(X,d_{X})$ to $(Y,d_{Y})$, we write $d_{\alpha}(x,y):=d_{Y}(\alpha(x),\alpha(y))$ to denote the distance between the images of two points $x,y\in X$ under this map.
$\mathcal{D}$ is a probabilistic embedding of $(X,d_{X})$ into $\mathcal{Y}$ if for each function $\alpha$ in the support of $\mathcal{D}$, $\alpha:X\rightarrow Y(\alpha)$ for some $(Y(\alpha),d_{Y(\alpha)})\in\mathcal{Y}$. This embedding has distortion $c$ if for all $u,v\in X$ we have:
@@
$\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq c\cdot d(u,v).$
Observe that every embedding in the support of $\mathcal{D}$ is expanding, [^1] Embeddings can also be defined without the expansion requirement, but we will stick to expanding embeddings in this paper. but pairwise distortion is bounded in expectation over the randomness in $\mathcal{D}$. If the support of $\mathcal{D}$ contains a single embedding, it is a deterministic embedding.
@@
If $\mathcal{D}$ is a probabilistic embedding from $(X,d_{X})$ into $\mathcal{Y}$ and $K\subseteq X$, define $\mathcal{D}_{X\setminus K}$ to be the distribution identical to $\mathcal{D}$ but with all functions $\alpha$ in the support restricted to the set $X\setminus K$ (i.e. we can sample from $\mathcal{D}_{X\setminus K}$ by sampling from $\mathcal{D}$ and restricting the domain of the function drawn).
Definition 1
@@
Note that in Definition 1, there is a deterministic subset $K$ whose distances to other nodes (and each other) we don’t care about. The randomness is only on the embedding for the remaining nodes in $X\setminus K$. One motivation for outlier embeddings is as a way to stop noise or adversarial data from preventing good embeddings, and we believe that making the outlier set itself deterministic is consistent with this motivation. In Section 5 we also discuss an application that naturally lends itself to finding a deterministic outlier set.
Nested embeddings
As mentioned in earlier, we extend the notion of nested embeddings from Chawla and Sheridan [19] to the probabilistic setting.
Definition 2(Probabilistic composition of nested embeddings, modified from [19])
@@
An $(\ell,g)$-nested composition of the assortment $(\mathcal{D}_{S},\set{\mathcal{D}_{K^{\prime}}}_{|K^{\prime}|\leq k+1})$, where $\ell$ is a constant and $g:[1,\infty)^{2}\times\mathbb{N}\rightarrow[1,\infty)$, is an algorithm that given the assortment returns a probabilistic embedding $\mathcal{D}:X\xhookrightarrow{}_{p}\mathcal{Y}$ such that:
@@
$\text{for all }u,v\in S,$ $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq\ell\cdot c_{S}\cdot d(u,v),$ (2)
$\text{and, for all }u,v\in X,$ $\mathop{\mathbb{E}}_{\alpha\sim\mathcal{D}}[d_{\alpha}(u,v)]$ $\leq g(c_{S},c_{X},k)\cdot d(u,v).$ (3)
@@
Definition 3
@@
  • Each node $u$ in the tree is associated with some value $\eta_{u}$, and $\eta_{v}=\beta\cdot\eta_{u}$ whenever $u$ is a child of vv. [^2] In this paper, we will use the term $\beta$-HSTs to refer to what other sources sometimes call exact $\beta$-HSTs. These sources typically define non-exact $\beta$-HSTs as only requiring $\eta_{v}\geq\beta\cdot\eta_{u}$. Further, for all leaves $v$, $\eta_{v}=1/\beta$.
  • For two leaves $x,y\in X$ with least common ancestor $z$, $d(x,y)=\eta_{z}$.
For simplicity, we will primarily focus on $2$-HSTs, and in referring to HSTs, we mean $2$-HSTs unless otherwise noted. Note that focusing on $\beta$-HSTs for a different constant $\beta$ at worst changes results by a constant factor.
@@
To achieve our main outlier embedding result (Theorem 1.1), we show in Section 3 that HSTs admit nested embeddings.
Theorem 2.1(HSTs have nested embeddings)